Solve.
-8(10x-4)+10(8x-4)=x-9

Answer by malakumar_kos@yahoo.com(315)


Solve.
-8(10x-4)+10(8x-4)=x-9
-8(10x-4)+10(8x-4) = x-9
=> -80x+32+80x-40 = x-9
=> -40+32 = x-9
=> -8+9 = x

=> x = 1